primefactorization of the number 20
Answer by jim_thompson5910(35256)

Solved by pluggable solver: Factor any number
20 is NOT a prime number: 20 = 2 * 2 * 5

Work Shown

20 is divisible by 2: 20 = 10 * 2.
10 is divisible by 2: 10 = 5 * 2.
5 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.